OpenStudy (yisrael012332):

What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5 HELP PLEASE.....

OpenStudy (yisrael012332):

@heretohelpalways

OpenStudy (yisrael012332):

@princeharryyy

OpenStudy (anonymous):

I am here

OpenStudy (princeharryyy):

:) @heretohelpalways

OpenStudy (yisrael012332):

can you help me please im completely lost

OpenStudy (anonymous):

find x in x−y=−3

OpenStudy (anonymous):

so x=y−3

OpenStudy (yisrael012332):

-3x?

OpenStudy (anonymous):

x=y−3

OpenStudy (anonymous):

then substitude (y−3) for x in x+3y=5

OpenStudy (anonymous):

so y−3+3y=5

OpenStudy (anonymous):

then simplify

OpenStudy (anonymous):

to get y=2

OpenStudy (anonymous):

then substitude 2 for y in x=y−3

OpenStudy (anonymous):

x=-1 and y=2

OpenStudy (yisrael012332):

but its only asking for the value of X so what do i put 2 or -1 even tho that's for Y.

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

so put -1

OpenStudy (anonymous):

damn your smart